OpenStudy (anonymous):

Please Help!! Express the complex number in trigonometric form. 3i

OpenStudy (ikram002p):

if u have complex number z=x+iy then u can write it as z= r(cos theta + i sin theta ) such that r=|z|

OpenStudy (jdoe0001):

3i => 0 + 3i x y thus 

OpenStudy (jdoe0001):

\(\large \bf z= r[cos(\theta)+isin(\theta)]\)

OpenStudy (anonymous):

im not understanding

OpenStudy (jdoe0001):

well.. you'd need to find keep in mind that the complex a + bi form correlates to the rectangular (x,y) coordinates from there, you get the magnitude or "r" or hypotenuse if you like and the angle

OpenStudy (jdoe0001):

what would be the "r" or magnitude of that vector? what's the angle?

OpenStudy (jdoe0001):

thus \(\large \bf {\color{blue}{ r}}[cos({\color{purple}{ \theta}})+isin({\color{purple}{ \theta}})]\)

OpenStudy (anonymous):

is r 3?

OpenStudy (jdoe0001):

x = 0 y = 3 \(\bf {\color{blue}{ r}}=\sqrt{a^2+b^2}\to \sqrt{x^2+y^2}\to \sqrt{0^2+3^2}\to 3\)

OpenStudy (ikram002p):

ok , r = length of z u can find it through \(r=\sqrt (x^2 +y^2)\) and as if u have complex number z=x+iy then u can write it as z= r(cos theta + i sin theta ) such that r=|z| then x= r cos theta y= r sin theta then find theta , r :D

OpenStudy (anonymous):

ok, so r is 3

OpenStudy (jdoe0001):

and the angle, should be quite obvious

OpenStudy (anonymous):

90

OpenStudy (anonymous):

can i ask for help on another one?

OpenStudy (jdoe0001):

yeap, thus \(\large \bf {\color{blue}{ 3}}[cos({\color{purple}{ 90^o}})+isin({\color{purple}{ 90^o}})]\)

OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 (sqrt 3)i

OpenStudy (jdoe0001):

same way really

OpenStudy (anonymous):

so r is 3sqrt3?

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} -3&+3\sqrt{3}\ i\\ \uparrow &\uparrow \\ x&y \end{array}\)

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ r}}=\sqrt{a^2+b^2}\to \sqrt{x^2+y^2}\)

OpenStudy (anonymous):

so 6

OpenStudy (anonymous):

how would i find the angle now?

OpenStudy (jdoe0001):

yeap so you'd find the "r" or magnitude and to find the angle... keep in mind that \(\bf tan(\theta)=\cfrac{y}{x}\to \cfrac{3\sqrt{3}\ i}{-3}\quad thus\quad {\color{purple}{ \theta}}=tan^{-1}\left(\cfrac{3\sqrt{3}\ i}{-3}\right)\)

OpenStudy (anonymous):

oh ok

OpenStudy (jdoe0001):

well... I should have removed the "i' from there =) \(\bf tan(\theta)=\cfrac{y}{x}\to \cfrac{3\sqrt{3}}{-3}\quad thus\quad {\color{purple}{ \theta}}=tan^{-1}\left(\cfrac{3\sqrt{3}}{-3}\right)\)

OpenStudy (anonymous):

so if i change that to cos and i sin would that be 6(cos (2pi/3) + i sin (2pi/3))

OpenStudy (anonymous):

wait no, it should be (5pi/6)

OpenStudy (jdoe0001):

yeap well.. keep in mind that tangent is negative in II and IV quadrants

OpenStudy (anonymous):

can you help me with one more? im half way through with it, but i don't know how to finish it

OpenStudy (anonymous):

Find the cube roots of 125(cos 288° + i sin 288°). cubed root(125*(cos(288)+ i sin (288)) cubed root(125)*cubed root(cos(288)+i sin(288)) 5*(cos(288)+ i sin(288))^1/3

OpenStudy (jdoe0001):

actaully should be the 1st one you listed \(\bf \cfrac{2\pi}{3} \quad and\quad \cfrac{5\pi}{3}\)

OpenStudy (anonymous):

oh ok

OpenStudy (jdoe0001):

is simpler posting anew... if I dunno someone else may, and thus more exposure and also we can reviise each other :)

OpenStudy (anonymous):

ok :)